OpenStudy (anonymous):

Use the quadratic formula to solve x^2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. A. {–12, 5} B. {–11.1, –2.9} C. {–5.6, –1.4} D. {–3.9, 12.1}

OpenStudy (jdoe0001):

\(\large \bf ax^2+bx+c\\\quad\\ \text{quadratic formula}\\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\)

OpenStudy (anonymous):

In other words.....

OpenStudy (jdoe0001):

\(\bf x^2 + 7x + 8 = 0\\ \quad \\ x= \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\implies x= \cfrac{ - (7) \pm \sqrt { (7)^2 -4(1)(8)}}{2(1)}\)

OpenStudy (anonymous):

Actual: x=−1.4384471871911697 and x=−5.561552812808831